OpenStudy (anonymous):

solve (7 , 52.50) m= 15/2 my final answer was -15x+2y=0 am i right?

OpenStudy (anonymous):

if there's any confusion it has to be in point-slope form.

OpenStudy (turingtest):

The version you have is in standard form. Point-slope form of a line passing through\[(x_1,y_1)=(7,52.5)\]looks like this:\[y-y_1=m(x-x_1)\]so in your case we have\[y-52.5=\frac{15}{2}(x-7)\]looks like you made a mistake trying to put it into standard form, but it's unnecessary anyway, so you don't even need to mess with it.

OpenStudy (turingtest):

...in standard form it would be\[2y-15x=-52.5\]so it looks like you lost the constant somehow, but you were kinda close.

OpenStudy (anonymous):

thank you for your help.